OpenStudy (joshoyen):

PLEASE HELP ASAP. WILL MEDAL AND FAN. Solve each equation. Check each solution. \[-\frac{ 4 }{ x+1 }=\frac{ 5 }{ 3x+1 }\]

OpenStudy (anonymous):

Any idea on how to start this?

OpenStudy (joshoyen):

No clue, really confused on this subject.

OpenStudy (anonymous):

Try to remember on how you moved parts/components across the equal sign. I can help guide you through this question!

OpenStudy (anonymous):

We can start by moving the denominators and multiplying it across the equal sign: Here I'll show you \[\frac{ -4 }{ (x+1) } = \frac{ 5 }{ 3x+1 }\] \[-4(3x+1) = 5(x+1) \]

OpenStudy (anonymous):

Can you solve the rest from here?

OpenStudy (joshoyen):

ohhh okay yeah i can do it now, thanks man

OpenStudy (anonymous):

No worries, glad to help!

OpenStudy (joshoyen):

@DelTaVsPi I have another question. for \[\frac{ x-4 }{ 3 }=\frac{ x-2 }{ 2 }\] would it be, 3(x-4) = x(x=2) ?

OpenStudy (anonymous):

The correct expression is: 2(x-4) = 3(x-2)

OpenStudy (sloppycanada):

Yup, precisely.

OpenStudy (anonymous):

From here you can continue, but your starting method is on point!

OpenStudy (joshoyen):

ohhh okay thanks guys!